# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 2.1.4 The distance from equilibrium
The three conditions on the distance function discussed above are, in fact, not enough to avoid several possible anomalies. This was noticed already in [5], where an additional demand was added somewhat “heuristically”: that the exponential ratio between the time-dependent coefficients of $\vec{v}_{2}$ and $\vec{v}_{3}$ not be compensated for by a distance function with an exponential weight in some directions relative to others. In other words, the metric should not grow exponentially faster in one direction than another, so that it compensates for the temporal exponential relaxation in different directions. This point was cleared up only in [60]. Using recent results on thermomajorization [72], the exact conditions for the existence of the Mpemba effect using any monotone as a distance function were found. Without getting into the details of thermomajorization, it was shown that a sufficient condition for having a Mpemba effect with any monotone as a distance function is
@@
It is important to note that the “thermomajorization Mpemba effect,” where the crossing between the distance function through relaxation must occur for any monotone, is a strong condition to impose. In most cases, a crossing of Kullback-Leibler divergences suffices to identify a Mpemba effect. The nonmonotonicity of $|a_{2}(T)|$ is then a sufficient demand for the existence of the effect: in the long-time limit, we can expand
@@
To illustrate the issues with measuring the distance to equilibrium via the average energy, Fig. 7 shows a numerical simulation of the average energy of an ensemble of particles relaxing in a tilted double-well potential that corresponds to the experiments in Ref. [6] (see Sec. 6.2). To understand the striking nonmonotonic relaxation, we note that the relaxation in a double-well potential proceeds in two stages (c.f. Sec. 2.1.8, below). In the first stage, particles typically relax to the nearest local potential well. In the second stage, the relative probabilities of being in each state (well) relax to their equilibrium values [75]. In the particular case shown in Fig. 7, there is an excess chance that particles relax to the lower-energy well, implying an average energy that is below the equilibrium value. The energy increases to that value in the second stage of relaxation. The fact that two states (here, at $t\approx 0.04$ s and at long times, as illustrated by the center and right inset distributions) have the same average energy implies a violation of one of the basic axioms for a metric that the distance $D[p,\pi]>0$ for all $p\neq\pi$. The nonmonotonicity of relaxation implies that two relaxation curves could cross without necessarily implying a Mpemba effect.
Figure 7: Numerical simulation of the average energy evolution relative to the equilibrium value for an ensemble of single-particle systems relaxing to equilibrium. The initial system is in equilibrium with a high-temperature bath, with nearly a uniform distribution (left inset, where the shaded pink distribution is nearly flat), and is highly different from the distribution in equilibrium with the bath (left insert, black curve). The average energy of the middle insert ($t\approx 0.04$ s) equals the value at long times (right insert). Numerical simulation courtesy of Siddharth Sane.
@@ 2.1.5 The inverse Mpemba effect
Figure 8: Inverse Markovian Mpemba effect in a three-state system. (a) Two identical systems, initiated at a cold temperature $T_{c}$ and a hot temperature $T_{h}$, are coupled to a hot bath with temperature $T_{b}>T_{h}>T_{c}$. The initially cold system has a very small component along the slow relaxation mode $\vec{v}_{2}$ (green dashed arrow). In contrast, the initially hot system has a much larger coefficient along the slow direction (even though it is at a smaller distance from the equilibrium). Hence, it decays more slowly than the initially cold system. Inset: Distance from equilibrium for both systems, as a function of time after quench. Initially, the cold system decays more quickly, and the inverse Markovian Mpemba effect occurs. Both distances ${D}[\vec{p}_{h}(t),\vec{\pi}(T_{b})]$ and ${D}[\vec{p}_{c}(t),\vec{\pi}(T_{b})]$ decrease with time. The initially colder system (blue dashed line) decays faster than the initially hotter one (red line), and the inverse Markovian Mpemba effect occurs. (b) Schematic of the asymmetric energy landscape $U(x)$ used to explore the inverse Mpemba effect, set asymmetrically within a one-dimensional box $[x_{\rm min},x_{\rm max}]$ with potential walls with large finite slope at the domain boundaries. Panel (a) is reproduced from [3] and panel (b) from [6].
Let us begin with an example of a three-state system. We set the energy levels of the three states to $E_{1}=0$, $E_{2}=0.1$ and $E_{3}=1$, while the heights of the barriers separating the wells are $B_{12}=2$, $B_{13}=1.01$ and $B_{23}=10$. The dynamics of the system is regulated by a rate matrix $\mathbf{W}$, whose components have an Arrhenius form, as outlined in Eq. (17). In contrast to the direct case, the inverse effect involves a heating protocol, meaning that the bath temperature is higher than both initial conditions. Setting, e.g., the bath temperature at $T_{b}=10$, we need to find a pair of temperatures $T_{c}<T_{h}<T_{b}=10$ for which one can observe a finite-time crossing of a distance measure between the probabilities $\vec{p}_{h,c}(t)$ and the target Boltzmann distribution $\vec{\pi}(T_{b})$.
@@ 2.1.6 The strong Mpemba effect
Figure 9: (a) Strong direct Mpemba effect, illustrated with a sketch highlighting the existence of a $T_{0}\neq T_{b}$ such that $a_{2}(T_{0},T_{b})=0$, which allows for exponentially faster cooling. (b) 1D potential for a Brownian colloidal particle that allows the observation of a manifestation of the strong Mpemba effect (see [6]), in which the distance measure ${D}$ exhibits an exponentially faster relaxation, as shown in panel (c). (d) The Boltzmann distribution in the well at infinite and bath temperature $T_{b}$, where the red (gray) area highlights how the total probability is partitioned between the two wells for both temperatures. The Mpemba effect is the strongest when the two red (gray) areas match for the initial (here sketched at $T_{h}\to\infty$) and the final (bath) temperature $T_{b}$: $\Pi_{L}(T_{h})\approx\Pi_{L}(T_{b})$. Panels b), c), and d) were adapted from Ref. [6].
@@
implying an exponentially faster convergence towards $\pi(\vec{x},T_{b})$, with a corresponding timescale given by $\lambda_{3}$ rather than by $\lambda_{2}$. The $a_{2}$ temperature dependence is exemplified in a sketch in Fig. 9a, which illustrates the strong, direct Mpemba effect scenario. In this example, there exists some $T_{b}<T_{c}<T_{h}$ such that $a_{2}(T_{h},T_{b})\equiv 0$. Note that $a_{2}(T_{0}=T_{b},T_{b})=0$, as well, since a system initially at equilibrium with the bath at temperature $T_{b}$ obeys $p(\vec{x},t)=\pi(\vec{x},T_{b})$ exactly, with all $a_{i>1}=0$. An analogous picture applies for a strong inverse effect for a heating protocol. The existence of a strong Mpemba effect implies the existence of the weak effect too, as $a_{2}(T_{0},T_{b})$ is a continuous function of $T_{0}$. Thus, if $a_{2}$ has two isolated zeroes in the strong-Mpemba-effect scenario, it must be a nonmonotonic function of initial temperature, implying the existence of the weak Mpemba effect in between the two zeros of $a_{2}$.
Figure 9 shows an experimental observation of the strong direct Mpemba effect in a colloidal setup, where a Brownian particle is diffusing in the 1D potential depicted in Fig. 9b. The $L_{1}$ distance measure for this system is observed to relax from an initial hot state (represented by the red line) and a cold state (represented by the blue line) towards an even colder bath temperature (indicated by the black line). Both the distance and time axes are shown on a logarithmic scale, which helps visualize the decay curve. The initially hot system exhibits a decay curve with a different slope compared to the relaxation from the cold initial condition, highlighting the exponentially faster cooling phenomenon [6]. Panel d) provides a sketch of the experimental findings [6] that are theoretically confirmed in Ref. [76]. It shows that the Mpemba effect is strongest when the equilibrium probability of the particle being in the left well at the bath temperature, $\Pi_{L}(T_{b})$, closely matches the probability of the particle being in the left well at the initial hot temperature, $\Pi_{L}(T_{h})$. This condition enables the system to attain global equilibrium by locally adjusting the probability distribution within each well rather than depending on transitions over the barrier. It is important to note that transitions between the wells are infrequent due to the small diffusion as well as the presence of a large and wide barrier.
@@
As a simple example of these ideas, we return to the discrete three-state system introduced in Fig. 4, which can be considered an effective representation of an overdamped particle in a three-well potential. The system can be described by a normalized probability vector $\vec{p}=(p_{1},p_{2},p_{3})$, with $\sum_{i}p_{i}=1$ and $p_{i}\geq 0$. States lie on the probability simplex highlighted with a blue triangle in Fig. 10, where the equilibrium locus, representing all the Boltzmann equilibrium vectors $\vec{\pi}(T)$ for $T\in[0,\infty]$ is highlighted as the blue to red curve lying on the probability simplex. Each component is related to the energies of the states by a Boltzmann factor $\pi_{i}(T)\propto e^{-E_{i}/k_{B}T}$, where $k_{B}$ is the Boltzmann constant. The locus is representative (without loss of generality) of a system with the lowest energy configuration in State 1. For $T=0$, it approaches the probability vector $(1,0,0)$ (blue side of the curve), while in the limit of infinite temperature the locus approaches the equipartition probability vector $(\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}})$ lying at the center of the simplex (red side of the curve). For a given bath temperature $T_{b}$ (black point marked by $T_{b}$ on the simplex), the set of all points $\vec{x}$ such that $\vec{x}\cdot\vec{u}_{2}(T_{b})=0$ is highlighted with an orange surface. Its intersection with the simplex provides all the probability vectors $\vec{p}$ such that $\vec{p}\cdot\vec{u}_{2}(T_{b})=0$. While the equilibrium locus intersects the $a_{2}=0$ plane at $\vec{\pi}(T_{b})$ by definition, the set of $\vec{p}$ resulting from the intersection with the simplex generally represents an out-of-equilibrium system. However, in the specific example highlighted in Fig. 10, one finds that this line intersects the equilibrium locus at some temperature $\vec{\pi}(T_{M})$ with $T_{M}\neq T_{b}$ (black dot marked by $T_{\rm{SM}}$) since the two extremes of the equilibrium locus (at $T=0$ and $T=\infty$) lie on the same side of the $a_{2}=0$ hyperplane: this provides a case for the strong Mpemba effect.
The hyperplane $a_{2}=0$ has codimension one, and hence, it separates the probability simplex into two disjoint sets. Therefore, assuming that the equilibrium locus is a continuous function of the temperature, namely that there are no phase transitions in the system, the parity of the number of times the equilibrium locus crosses $a_{2}=0$ depends solely on the position of its boundaries at $T=0$ and $T=\infty$. Given a certain $T_{b}$, a sufficient condition for the strong Mpemba effect is that $\textrm{sign}[a_{2}(T_{0}=0,T_{b})]=\textrm{sign}[a_{2}(T_{0}=+\infty,T_{b})]$. With this motivation, we can define the parity of $\mathcal{I}_{M}$ in terms of the Heaviside step function $\theta$ as
@@
Mathematically, a situation in which $a_{2}(T_{0},T_{b})=0$ without a sign change (i.e., $\partial_{T_{0}}a_{2}(T_{0},T_{b})=0$) can occur. This scenario differs from the parity characterization outlined above but requires fine-tuning of the system parameters; a small perturbation would generally destroy the needed degeneracy. Nevertheless, while in these cases, $\mathcal{P}(\mathcal{I}_{M})$ does not correspond to the exact parity, it still serves as a lower bound to the number of crossings. We note that a discontinuity in $a_{2}(T_{0},T_{b})$ with respect to a change in the bath temperature can appear if the second and third eigenvalues become equal at some $T_{b}$ (i.e., if $\lambda_{2}=\lambda_{3}$). This case is discussed in Sec. 9.3 and, in fact, generates the transition into the strong inverse effect in the Ising antiferromagnet at low values of the magnetic field [5, 70].
In Sec. 7.1.3, it is shown that the strong effect turned out to be extremely useful in identifying the Mpemba effect in numerical simulations of large systems, where the number of microstates is way too large to enable reasonable sampling of the evolving distribution. In such cases, Monte-Carlo simulations can be used to identify the strong Mpemba (direct and inverse) effects, but no similar method is known to identify the weak effects.
@@ 2.1.7 Mpemba effect in boundary and weak coupling setups
The theoretical models exhibiting the Mpemba effect discussed so far assumed that all relevant degrees of freedom (DoF) are directly coupled to the thermal bath. However, in most experiments demonstrating the Mpemba effect—the colloidal system [6, 64] being the only exception as it has a single DoF—the system is coupled to the heat bath only through its boundaries, so that the heat is exchanged through a limited portion of the system (e.g., its surface). One might then worry that instances of the Mpemba effect studied so far require that all the DoF be fully coupled to the thermal bath and that the models of the Mpemba effect discussed above may not be relevant to the observations of anomalous relaxation in other systems, which are spatially extended. Intuitively, one would expect the effect to be more visible as the coupling with the thermal bath strengthens, while a weak coupling, on the other hand, should make it harder (if not impossible) to observe. A Markovian framework for a thermal quench through boundary coupling was developed in Ref. [70], and it was used to study the influence of the coupling strength on the presence of a Mpemba effect.
Let us consider a system that is coupled to a thermal bath only through its boundaries. In other words, heat can be transferred between the system and the bath only through those DoF sitting on the boundaries. All other transitions between microstates are “bulk transitions”, and as no energy is exchanged with the bath in those transitions. Therefore, due to the conservation of total energy, they can only happen between states with the same energy. These bulk transitions serve as a self-ergodizing (SE) mechanism, whereas the boundary transitions (BT) couple the system to the bath and enable transitions between different energy shells. This structure can be modeled by
@@
Here $\mathbf{W}^{\rm SE}$ and $\mathbf{W}^{\rm BC}$ are normalized rate matrices corresponding to the self-ergodizing and boundary coupling transitions, respectively, and $\Gamma^{\{\rm SE,\rm BC\}}$ are coupling constants modulating the rates amplitude. The ratio $C=\Gamma^{\rm BC}/\Gamma^{\rm SE}$ determines the coupling strength: in the limit $C\ll 1$, boundary flips— in which energy can be exchanged with the bath— occur rarely compared to ergodization flips, so the system ergodizes quickly after each energy exchange with the bath. In the $C\gg 1$ limit, the boundaries exchange heat much faster than the ergodization rate. Thus, the boundaries are in thermal equilibrium with the bath, and the diffusion of energy into the system sets the timescale for the relaxation. The specific normalization chosen for $\mathbf{W}^{\{\rm BC,\rm SE\}}$ changes the value of $C$, but not its limiting cases. As an example, one can choose to normalize with respect to the maximum rate so that $\max\left(W_{i\!j}^{\rm SE}\right)=\max\left(W_{i\!j}^{\rm BC}\right)\equiv 1$.
By construction, $\mathbf{W}^{\rm SE}$ is a rate matrix that contains only transitions between states that have the same energy. Generically, it is a reducible matrix with a zero eigenvalue degeneracy equal to the number of different energy macrostates. As for $\mathbf{W}^{\rm BC}$, it is generically a sparse matrix, as most transitions involve more than the boundary spins. The interaction of the system with a thermal bath at temperature $T_{b}$ satisfies detailed balance; therefore, $W^{\rm BC}_{ji}=W^{\rm BC}_{i\!j}e^{-(E_{j}-E_{i})/T_{b}}$, where the $E_{\{i,j\}}$ are the energies corresponding to the two states.
@@
Let us demonstrate the above construction with a specific example (see the sketch of Fig. 11a) consisting of a one-dimensional ring of $N$ Ising spins with nearest-neighbor antiferromagnetic interactions. Each spin $\{\sigma_{s}\}_{s=1\dots N}$ in the chain can either be in an up or down (±$\pm 1$) state, giving a total of $M=2^{N}$ different microstates, identified by the $N$-dimensional vector $\vec{\sigma}$. The energy of a microstate $\vec{\sigma}$ is provided by the Hamiltonian function
@@
$W^{\rm BC}_{i\!j}=\frac{\delta_{\sigma_{1}^{i},-\sigma_{1}^{j}}\prod_{s>1}\delta_{\sigma_{s}^{i},\sigma_{s}^{j}}}{1+e^{\beta_{b}\left(\mathcal{H}\left(\vec{\sigma}^{i}\right)-\mathcal{H}\left(\vec{\sigma}^{j}\right)\right)}}$ (45)

### v3

@@ 2.1.4 The distance from equilibrium
The three conditions on the distance function discussed above are, in fact, not enough to avoid several possible anomalies. This was noticed already in [5], where an additional demand was added somewhat “heuristically”: that the exponential ratio between the time-dependent coefficients of $\vec{v}_{2}$ and $\vec{v}_{3}$ not be compensated for by a distance function with an exponential weight in some directions relative to others. In other words, the metric should not grow exponentially faster in one direction than another, so that it compensates for the temporal exponential relaxation in different directions. This point was cleared up only in [60], where the exact conditions for the existence of the Mpemba effect using any monotone as a distance function were found, using recent results on thermomajorization [72]. In particular, a sufficient condition for having a Mpemba effect with any monotone as a distance function is
@@
It is important to note that the “thermomajorization Mpemba effect,” where the crossing between the distance function through relaxation must occur for any monotone, is a strong condition to impose. In most cases, a crossing of Kullback-Leibler divergences suffices to identify a Mpemba effect. The nonmonotonicity of $|a_{2}(T)|$ is then a sufficient condition for the existence of the effect: in the long-time limit, we can expand
@@
To illustrate the issues with measuring the distance to equilibrium via the average energy, Fig. 7a shows a numerical simulation of the average energy of an ensemble of particles relaxing in a tilted double-well potential that corresponds to the experiments in Ref. [6]; see Sec. 6.2. To understand the striking nonmonotonic relaxation, we note that the relaxation in a double-well potential proceeds in two stages (see also Sec. 2.1.8, below). In the first stage, particles typically relax to the nearest local potential well. In the second stage, the relative probabilities of being in each state (well) relax to their equilibrium values [75]. In the particular case shown in Fig. 7, more particles relax to the lower-energy well, implying an average energy that is below the equilibrium value. The energy increases to that value in the second stage of relaxation. The fact that two states (here, at $t\approx 0.04$ s and at long times, as illustrated by the center and right inset distributions) have the same average energy violates one of the basic axioms for a metric, that $D[p,\pi]>0$ for all $p\neq\pi$. The nonmonotonicity of relaxation implies that two relaxation curves could cross without necessarily implying a Mpemba effect. Indeed, Fig. 7b shows the “standard” L1 distance plot for the same simulation and confirms that there is no Mpemba effect for this case (no crossing of hot and cold curves), even though the average energy curves do cross.
Figure 7: Average energy is not always a reliable indicator of the Mpemba effect. (a) Numerical simulation of the average energy evolution relative to the equilibrium value for an ensemble of single-particle systems relaxing to equilibrium from hot (red curve) and cold (blue curve) initial temperatures, along with a control (black curve) that is initially in equilibrium with the bath. The initial “hot” system has a nearly uniform distribution (left inset, where the shaded pink distribution is nearly flat), and differs significantly from the distribution in equilibrium with the bath (right insert). The average energy of the middle, “cold” temperature initial state (blue curve) equals that of the hot initial temperature (red curve) at $t\approx 0.04$ s. (b) $L_{1}$ distance for the same numerical simulations show no crossing of hot (red) and cold (blue) curves and thus no Mpemba effect. Numerical simulations courtesy of Siddharth Sane.
@@ 2.1.5 The inverse Mpemba effect
Figure 8: Inverse Markovian Mpemba effect in a three-state system. (a) Two identical systems, initiated at a cold temperature $T_{c}$ and a hot temperature $T_{h}$, are coupled to a hot bath with temperature $T_{b}>T_{h}>T_{c}$. The initially cold system has a very small component along the slow relaxation mode $\vec{v}_{2}$ (green dashed arrow). In contrast, the initially hot system has a much larger coefficient along the slow direction (even though it is at a smaller distance from the equilibrium). Hence, it decays more slowly than the initially cold system. Inset: Distance from equilibrium for both systems, as a function of time after quench. Initially, the cold system decays more quickly, and the inverse Markovian Mpemba effect occurs. Both distances ${D}[\vec{p}_{h}(t),\vec{\pi}(T_{b})]$ and ${D}[\vec{p}_{c}(t),\vec{\pi}(T_{b})]$ decrease with time. The initially colder system (blue dashed line) decays faster than the initially hotter one (red line), and the inverse Markovian Mpemba effect occurs. (b) Schematic of the asymmetric energy landscape $U(x)$ used to explore the inverse Mpemba effect, set asymmetrically within a one-dimensional box $[x_{\rm min},x_{\rm max}]$ with potential walls with large finite slope at the domain boundaries. Source: Reprinted with permission from [3] and [6] respectively.
Let us begin with an example of a three-state system. We set the energy levels of the three states to $E_{1}=0$, $E_{2}=0.1$ and $E_{3}=1$, while the heights of the barriers separating the wells are $B_{12}=2$, $B_{13}=1.01$ and $B_{23}=10$. The dynamics of the system is regulated by a rate matrix $\mathbf{W}$, whose components have an Arrhenius form, as outlined in Eq. (17). In contrast to the direct case, the inverse effect involves a heating protocol, meaning that the bath temperature is higher than both initial conditions. Setting, for example, the bath temperature at $T_{b}=10$, we need to find a pair of temperatures $T_{c}<T_{h}<T_{b}=10$ for which one can observe a finite-time crossing of a distance measure between the probabilities $\vec{p}_{h,c}(t)$ and the target Boltzmann distribution $\vec{\pi}(T_{b})$.
@@ 2.1.6 The strong Mpemba effect
Figure 9: (a) Strong direct Mpemba effect, illustrated with a sketch highlighting the existence of a $T_{0}\neq T_{b}$ such that $a_{2}(T_{0},T_{b})=0$, which allows for exponentially faster cooling. (b) 1D potential for a Brownian colloidal particle that allows the observation of a manifestation of the strong Mpemba effect (see [6]), in which the distance measure ${D}$ exhibits an exponentially faster relaxation, as shown in panel (c). (d) The Boltzmann distribution in the well at infinite and bath temperature $T_{b}$, where the red (gray) area highlights how the total probability is partitioned between the two wells for both temperatures. The Mpemba effect is the strongest when the two red (gray) areas match for the initial (here sketched at $T_{h}\to\infty$) and the final (bath) temperature $T_{b}$: $\Pi_{L}(T_{h})\approx\Pi_{L}(T_{b})$. Source: Panels (b-d) are reprinted with permission from [6].
@@
implying an exponentially faster convergence towards $\pi(\vec{x},T_{b})$, with a corresponding rate given by $\lambda_{3}$ rather than by $\lambda_{2}$. The $a_{2}$ temperature dependence is exemplified in a sketch in Fig. 9a, which illustrates the strong, direct Mpemba effect scenario. In this example, there exists some $T_{b}<T_{c}<T_{h}$ such that $a_{2}(T_{h},T_{b})\equiv 0$. Note that $a_{2}(T_{0}=T_{b},T_{b})=0$, as well, since a system initially at equilibrium with the bath at temperature $T_{b}$ obeys $p(\vec{x},t)=\pi(\vec{x},T_{b})$ exactly, with all $a_{i>1}=0$. An analogous picture applies for a strong inverse effect for a heating protocol. The existence of a strong Mpemba effect implies the existence of the weak effect too, as $a_{2}(T_{0},T_{b})$ is a continuous function of $T_{0}$. Thus, if $a_{2}$ has two isolated zeroes in the strong-Mpemba-effect scenario, it must be a nonmonotonic function of initial temperature, implying the existence of the weak Mpemba effect in between the two zeros of $a_{2}$.
Figure 9 shows an experimental observation of the strong direct Mpemba effect in a colloidal setup, where a Brownian particle is diffusing in the 1D potential depicted in Fig. 9b. The $L_{1}$ distance measure for this system is observed to relax from an initial hot state (represented by the red line) and a cold state (represented by the blue line) towards an even colder bath temperature (indicated by the black line). Both the distance and time axes are shown on a logarithmic scale, which helps visualize the decay curve. The initially hot system exhibits a decay curve with a different slope compared to the relaxation from the cold initial condition, highlighting the exponentially faster cooling phenomenon [6]. Panel (d) provides a sketch of the experimental findings [6], which are theoretically confirmed in Ref. [76]. It shows that the Mpemba effect is strongest when the equilibrium probability of the particle being in the left well at the bath temperature, $\Pi_{L}(T_{b})$, closely matches the probability of the particle being in the left well at the initial hot temperature, $\Pi_{L}(T_{h})$. This condition enables the system to attain global equilibrium by locally adjusting the probability distribution within each well rather than depending on transitions over the barrier. It is important to note that transitions between the wells are infrequent because of the small diffusion coefficient and the large, wide barrier.
@@
As a simple example of these ideas, we return to the discrete three-state system introduced in Fig. 4, which can be considered an effective representation of an overdamped particle in a three-well potential. The system can be described by a normalized probability vector $\vec{p}=(p_{1},p_{2},p_{3})$, with $\sum_{i}p_{i}=1$ and $p_{i}\geq 0$. States lie on the probability simplex highlighted with a blue triangle in Fig. 10, where the equilibrium locus, representing all the Boltzmann equilibrium vectors $\vec{\pi}(T)$ for $T\in[0,\infty]$ is highlighted as the blue to red curve lying on the probability simplex. Each component is related to the energies of the states by a Boltzmann factor $\pi_{i}(T)\propto e^{-E_{i}/k_{B}T}$. The locus is representative (without loss of generality) of a system with the lowest energy configuration in State 1. For $T=0$, it approaches the probability vector $(1,0,0)$ (blue side of the curve), while in the limit of infinite temperature the locus approaches the equipartition probability vector $(\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}})$ lying at the center of the simplex (red side of the curve). For a given bath temperature $T_{b}$ (black point marked by $T_{b}$ on the simplex), the set of all points $\vec{x}$ such that $\vec{x}\cdot\vec{u}_{2}(T_{b})=0$ is highlighted with an orange surface. Its intersection with the simplex provides all the probability vectors $\vec{p}$ such that $\vec{p}\cdot\vec{u}_{2}(T_{b})=0$. While the equilibrium locus intersects the $a_{2}=0$ plane at $\vec{\pi}(T_{b})$ by definition, the set of $\vec{p}$ resulting from the intersection with the simplex generally represents an out-of-equilibrium system. However, in the specific example highlighted in Fig. 10, this line intersects the equilibrium locus at some temperature $\vec{\pi}(T_{M})$ with $T_{M}\neq T_{b}$ (black dot marked by $T_{\rm{SM}}$) since the two extremes of the equilibrium locus (at $T=0$ and $T=\infty$) lie on the same side of the $a_{2}=0$ hyperplane: this provides a case for the strong Mpemba effect.
Because the hyperplane $a_{2}=0$ has codimension one, it separates the probability simplex into two disjoint sets. Therefore, assuming that the equilibrium locus is a continuous function of the temperature, namely that there are no phase transitions in the system, the parity of the number of times the equilibrium locus crosses $a_{2}=0$ depends solely on the position of its boundaries at $T=0$ and $T=\infty$. Given a certain $T_{b}$, a sufficient condition for the strong Mpemba effect is that $\textrm{sign}[a_{2}(T_{0}=0,T_{b})]=\textrm{sign}[a_{2}(T_{0}=+\infty,T_{b})]$. With this motivation, we can define the parity of $\mathcal{I}_{M}$ in terms of the Heaviside step function $\theta$ as
@@
Mathematically, a situation in which $a_{2}(T_{0},T_{b})=0$ without a sign change (i.e., $\partial_{T_{0}}a_{2}(T_{0},T_{b})=0$) can occur. This scenario differs from the parity characterization outlined above but requires fine-tuning of the system parameters: a small perturbation would generally destroy the needed degeneracy. Nevertheless, while in these cases, $\mathcal{P}(\mathcal{I}_{M})$ does not correspond to the exact parity, it still serves as a lower bound to the number of crossings. We note that a discontinuity in $a_{2}(T_{0},T_{b})$ with respect to a change in the bath temperature can appear if the second and third eigenvalues become equal at some $T_{b}$ (i.e., if $\lambda_{2}=\lambda_{3}$). This case is discussed in Sec. 9.3 and, in fact, generates the transition into the strong inverse effect in the Ising antiferromagnet at low values of the magnetic field [5, 70].
In Sec. 7.1.3, it is shown that the strong effect turned out to be extremely useful in identifying the Mpemba effect in numerical simulations of large systems, where the number of microstates is so large that a reasonable sampling of the evolving distribution is impossible. In such cases, Monte-Carlo simulations can be used to identify the strong Mpemba (direct and inverse) effects, but no similar method is known to identify the weak effects.
@@ 2.1.7 Mpemba effect in boundary and weak coupling setups
The theoretical models exhibiting the Mpemba effect discussed so far assumed that all relevant degrees of freedom (DoF) are directly coupled to the thermal bath. However, in most experiments demonstrating the Mpemba effect—the colloidal system [6, 64] being the only exception as it has a single DoF—the system is coupled to the heat bath only through its boundaries, so that the heat is exchanged through a limited portion of the system (e.g., its surface). One might then worry that instances of the Mpemba effect studied so far require that all the DoF be fully coupled to the thermal bath and that the models of the Mpemba effect discussed above may not be relevant to the observations of anomalous relaxation in other systems, which are spatially extended. Intuitively, one would expect the effect to be more visible as the coupling with the thermal bath strengthens, while a weak coupling would make it harder to observe, if not impossible. A Markovian framework for a thermal quench through boundary coupling was developed in Ref. [70] and used to study the influence of the coupling strength on the presence of a Mpemba effect.
Let us consider a system that is coupled to a thermal bath only through its boundaries. In other words, heat can be transferred between the system and the bath only through those DoF sitting on the boundaries. All other transitions between microstates are “bulk transitions,” and no energy is exchanged with the bath in those transitions. From conservation of total energy, they can happen only between states with the same energy. These bulk transitions serve as a self-ergodizing (SE) mechanism, whereas the boundary transitions (BT) couple the system to the bath and enable transitions between different energy shells. This structure can be modeled by
@@
Here $\mathbf{W}^{\rm SE}$ and $\mathbf{W}^{\rm BC}$ are normalized rate matrices corresponding to the self-ergodizing and boundary coupling transitions, respectively, and $\Gamma^{\{\rm SE,\rm BC\}}$ are coupling constants modulating the rates amplitude. The ratio $C=\Gamma^{\rm BC}/\Gamma^{\rm SE}$ determines the coupling strength: in the limit $C\ll 1$, boundary flips—in which energy can be exchanged with the bath—occur rarely compared to ergodization flips, and the system ergodizes quickly after each energy exchange with the bath. In the $C\gg 1$ limit, the boundaries exchange heat much faster than the ergodization rate. Thus, the boundaries are in thermal equilibrium with the bath, and the diffusion of energy into the system sets the timescale for the relaxation. The specific normalization chosen for $\mathbf{W}^{\{\rm BC,\rm SE\}}$ changes the value of $C$, but not its limiting cases. As an example, one can choose to normalize with respect to the maximum rate so that $\max\left(W_{i\!j}^{\rm SE}\right)=\max\left(W_{i\!j}^{\rm BC}\right)\equiv 1$.
By construction, $\mathbf{W}^{\rm SE}$ is a rate matrix that contains only transitions between states that have the same energy. Generically, it is a reducible matrix with a zero-eigenvalue degeneracy equal to the number of different energy macrostates. As for $\mathbf{W}^{\rm BC}$, it is generically a sparse matrix, as most transitions involve more than the boundary spins. The interaction of the system with a thermal bath at temperature $T_{b}$ satisfies detailed balance; therefore, $W^{\rm BC}_{ji}=W^{\rm BC}_{i\!j}e^{-(E_{j}-E_{i})/k_{B}T_{b}}$, where the $E_{\{i,j\}}$ are the energies corresponding to the two states.
@@
As sketched in Fig. 11a, we illustrate the above construction using a one-dimensional ring of $N$ Ising spins with nearest-neighbor antiferromagnetic interactions. Each spin $\{\sigma_{s}\}_{s=1\dots N}$ in the chain can either be in an up or down (±$\pm 1$) state, giving a total of $M=2^{N}$ different microstates, identified by the $N$-dimensional vector $\vec{\sigma}$. The energy of a microstate $\vec{\sigma}$ is provided by the Hamiltonian function
@@
$W^{\rm BC}_{i\!j}=\frac{\delta_{\sigma_{1}^{i},-\sigma_{1}^{j}}\prod_{s>1}\delta_{\sigma_{s}^{i},\sigma_{s}^{j}}}{1+e^{\beta_{b}\left(\mathcal{H}\left(\vec{\sigma}^{i}\right)-\mathcal{H}\left(\vec{\sigma}^{j}\right)\right)}},$ (45)
@@
To model bulk transitions, we use rates that decay exponentially as $2^{-d_{i\!j}}$, where $d_{i\!j}=\sum_{s}\delta_{\sigma_{s}^{i},-\sigma_{s}^{j}}$ is the Hamming distance [81] that counts the number of spins that must be flipped between the two configurations, as suggested by decimation-like procedures operated on Markov jump processes [78]. We therefore formalize bulk transitions between two states $i\neq j$ as
The full transition matrix for the model is finally built as a linear combination of the two rate matrices, as in Eq. (43). With this construction, let us consider the persistence of the Mpemba effect in a boundary-coupling setup. In Fig. 11b, we plot the Mpemba phase diagram for $N=10$ (implying $M=1024$ microstates) as a function of bath temperature $\beta_{b}^{-1}$ and external magnetic field $H$. The resemblance with that of the model studied in [5] is striking, especially considering the profound dissimilarities between the two models: the all-to-all connectivity of a mean-field description makes it relevant for high-dimensional setups, while the model presented here addresses a one-dimensional spin chain. Additionally, in this model, we imposed a partial coupling with the thermal bath, while in the standard mean-field Ising antiferromagnet, every spin exchanges heat with the thermal bath. The similarity is not limited only to the examples where the Mpemba effect can be observed but extends to the many specific flavors of the Mpemba effect; see Fig. 12 for a detailed analysis. In general, we observe both direct/inverse, weak/strong and single/multiple effects, which are encoded by the different colors in the figure: here, boundary coupling merely shifts phase-diagram borders, and all flavors of the Mpemba effect can be observed. This hints that the antiferromagnetic character lies at the origin of the Mpemba effect, as it is the only feature shared among these very different models. The horizontal line at $H=2$ is determined by the phase transition that characterizes the system at $T_{b}=0$: for $0<H\leq 2$ (alternating spins ground state), only single Mpemba effects can be observed, while for $H>2$ (aligned spins ground state) double effects occur. For $0<H\leq 2$, a strong inverse Mpemba effect effect exists for $T_{b}\geq T_{b}^{*}=1/\beta^{*}$ (highlighted with a red arrow). The reason that no Mpemba effect exists for $T<T_{b}^{*}$, for small external magnetic fields, traces back to a phase transition of the relaxation dynamics [82]; see the discussion of applications in Sec. 8.